it took the pilot an hour and a half to make a flight of 240 miles when flying a against headwind. The return trip took an hour and twelve minutes(the wind had not shifted nor changed its speed.) What it the speed of the wind?
Answer by mananth(16946)

Plane speed speed= x mph
Wind speed speed= y mph
with wind speed= (x+y)
against wind speed (x-y)
d= 240
240 / (x+y) = 1.2
divide by 1.2
200 / (x+y) = 1
(x+y) = 200 ............1
d= = 240
27/(x-y) = 1.20
divide by 1.20
240 /(x-y) = 1.5
x - y = 160 .............2
add up (1) & (2)
2 x = 360
/ 2
x= 180 mph speed of Plane speed in still air
plug value of x in (1)
we get y= 20 mph speed of Wind speed
m.ananth@hotmail.ca